Solve this problem involving joint variations
The area A of a triangle varies jointly as the base b and the height h. A triangle with a base of 8cm and a height of 9 cm has an are of 36 square centimeters. Find the area A when the base b is 10 cm and the height h is 7cm.
Step 1: Write the correct equation
Step 2: Use information given the problem to find the constant of variation k.
Step 3: Rewrite the equation in step 1 and substitute the value of k found in step 2.
Step 4: Substitute to the given equation the values given in the problem. When solving word problems, remember to include units in your final answer.

Answer by ewatrrr(24785)


Hi
Step 1: Write the correct equation
   A= kbh
Step 2: Use information given the problem to find the constant of variation k.
   36cm 2 = k8cm*9cm
    k = 1/2
Step 3: Rewrite the equation in step 1 and substitute the value of k found in step 2.
   A = (1/2)bh
Step 4: Substitute to the given equation the values given in the problem. 
  A = (1/2)(10cm)(7cm) = 35 cm^2 

Wish You the Best in your Studies.